# Effective Learning with Node Perturbation in Multi-Layer Neural Networks

### v5

@@ Abstract
Backpropagation (BP) remains the dominant and most successful method for training parameters of deep neural network models. However, BP relies on two computationally distinct phases, does not provide a satisfactory explanation of biological learning, and can be challenging to apply for training of networks with discontinuities or noisy node dynamics. By comparison, node perturbation (NP) proposes learning by the injection of noise into network activations, and subsequent measurement of the induced loss change. NP relies on two forward (inference) passes, does not make use of network derivatives, and has been proposed as a model for learning in biological systems. However, standard NP is highly data inefficient and unstable due to its unguided noise-based search process. In this work, we investigate different formulations of NP and relate it to the concept of directional derivatives as well as combining it with a decorrelating mechanism for layer-wise inputs. We find that a closer alignment with directional derivatives together with input decorrelation at every layer strongly enhances performance of NP learning with large improvements in parameter convergence and much higher performance on the test data, approaching that of BP. Furthermore, our novel formulation allows for application to noisy systems in which the noise process itself is inaccessible.
@@ 1 Introduction
Backpropagation (BP) is the workhorse of modern artificial intelligence. It provides an efficient way of performing multi-layer credit assignment, given a differentiable neural network architecture and loss function (Linnainmaa, 1970). Despite BP’s successes, it requires an auto-differentiation framework for the backward assignment of credit, introducing a distinction between a forward, or inference phase, and a backward, or learning phase, increasing algorithmic complexity and impeding implementation in (neuromorphic) hardware (Kaspar et al., 2021; Zenke & Neftci, 2021). Furthermore, BP has long been criticized for its lack of biological detail and plausibility (Grossberg, 1987; Crick, 1989; Lillicrap et al., 2016), with significant concerns again being the two separate forward and backward phases, but also its reliance on gradient propagation and the non-locality of the information required for credit assignment.
Alternative algorithms have been put forth over the years, though their inability to scale and difficulty in achieving levels of performance comparable to BP have held back their use. Many focus either on greedy local learning rules or on random feedback matrices and projections, which are deemed more plausible than symmetric feedback weights (Dellaferrera & Kreiman, 2022; Frenkel et al., 2021; Kohan et al., 2018; Hinton, 2022)
The category of algorithms we focus on in this work are perturbation algorithms, such as node perturbation (NP) (Dembo & Kailath, 1990; Cauwenberghs, 1992), which require no feedback connectivity at all. In NP, node activations are perturbed by a small amount of random noise. Weights are then updated to produce the perturbed activations in proportion to the degree by which the perturbation improved performance. This method requires a measure of the loss function being optimized on a network both before and after the inclusion of noise. Such an approach is appealing because it leverages the same forward pass twice, rather than relying on two computationally distinct phases. It also does not require non-local information other than the global performance signal.
@@
In this work, we put forward three contributions. First, we reframe the process of node perturbation together with the subsequent measurement of the output loss change in terms of directional derivatives. This provides a more solid theoretical foundation for node perturbation and, consequently, a different update rule, which we refer to as iterative node perturbation. Directional derivatives have been related to NP before in the context of forward gradient learning (Baydin et al., 2022; Ren et al., 2022). This approach is different from ours in that the perturbations are used to estimate gradients, but node-perturbed forward passes are not performed network-wide, making these approaches unsuitable for use in true noisy systems.
Second, we introduce an approximation to iterative node perturbation, referred to as activity-based node perturbation, which is more efficient and has the additional advantage that it can be implemented in noisy systems such as imprecise hardware implementations (Gokmen, 2021) or biological systems (Faisal et al., 2008), where the noise itself is not directly measurable.
Third, we propose to use a decorrelation method, first described by Ahmad et al. (2023), to debias layer-wise activations and thereby achieve faster learning. Because NP-style methods directly correlate perturbations of unit activities with changes in a reward signal, decorrelation of these unit activities helps to eliminate confounding effects, making credit assignment more straightforward. In addition, as demonstrated by Hiratani et al. (2022), correlations in the input data lead to more bias in NP’s updates by increasing their covariance. By combining decorrelation with the different NP methods, we find that it is possible to achieve orders of magnitude increase in model convergence speed, with performance levels rivaling networks trained by BP in certain contexts.
@@ 2.1 Node perturbation and its formulations
2.1.1 Traditional node perturbation
In the most common formulation of NP, noise is injected into each layer’s pre-activations and weights are updated in the direction of the noise if the loss improves and in the opposite direction if it worsens. Two forward passes are required: one clean and one noise-perturbed. During the noisy pass, noise is injected into the pre-activation of each layer to yield a perturbed output
Having defined a clean and noise-perturbed network-pass, we can measure a loss differential for a NP-based update. Supposing that the loss $\mathcal{L}$ is measured using the network outputs, the loss difference between the clean and noisy network is given by $\delta{\mathcal{L}}=\mathcal{L}(\mathbf{\tilde{x}}_{L})-\mathcal{L}(\mathbf{x}_{L})$, where $\delta{\mathcal{L}}$ is a scalar measure of the difference in loss induced by the addition of noise to the network. Given this loss difference and the network’s perturbed and unperturbed outputs, we compute a layer-wise learning signal by replacing the gradient $\mathbf{g}_{l}$ in Eq. 1 with a term consisting of the loss difference and a normalized noise vector:
$\Delta\mathbf{W}_{l}^{\textrm{NP}}=\delta{\mathcal{L}}\frac{{\bf\it\epsilon}_{l}}{\sigma^{2}}\mathbf{x}_{l-1}^{\top}\,.$ (3)
@@ 2.1.2 Iterative node perturbation
Though mathematically simple, the traditional NP approach described above has a number of drawbacks. One of the biggest drawbacks is that each layer’s noise impact upon the loss is confounded by additional noise in previous and following layers which is unaccounted for. Appendix A describes how correlations arise between layers. Furthermore, a precise relationship between the traditional NP formulation and gradient descent is missing.
In the following, we develop a more principled approach to node perturbation. Our goal is to determine the gradient of the loss with respect to the pre-activations in a layer $l$ by use of perturbations. To this end, we consider the partial derivative of the loss with respect to the pre-activation $\mathbf{a}^{i}_{l}$ of unit $i$ in layer $l$ for all $i$. We define a perturbed state as $\mathbf{\tilde{x}}_{k}(h)=f\left(\mathbf{\tilde{a}}_{k}+h\mathbf{m}_{k}\right)$ with $h$ an arbitrary scalar and binary vectors $\mathbf{m}_{k}=\mathbf{e}_{i}$ if $k=l$ with $\mathbf{e}_{i}$ a standard unit vector and $\mathbf{m}_{k}=\mathbf{0}$ otherwise.
We may now define the partial derivatives as
@@
An alternative approach is to define perturbations in terms of directional derivatives. Directional derivatives measure the derivative of a function based upon an arbitrary vector direction in its dependent variables. However, this can only be accomplished for a set of dependent variables which are individually independent of one-another. Thus, we cannot compute such a measure across an entire network. We can, however, measure the directional derivative with respect to a specific layer via a perturbation given by
@@ 2.1.3 Activity-based node perturbation
Though theoretically grounded, INP incurs significant overhead compared to the direct noise-based update method. Specifically, in order to precisely compute directional derivatives, noise must be applied in a purely layer-wise fashion, requiring a high degree of control over the noise injection. This is less biologically plausible and functionally more challenging implementation with less potential for scaling. Furthermore, the number of forward passes required to compute such an update for a whole network for INP is then the number of layers plus one, $L+1$, potentially many times larger than the two passes required for NP.
To balance learning speed and computational cost we instead approximate the directional derivative across the whole network simultaneously. This involves assuming that all node activations in our network are independent and treating the entire network as if it were a single layer. This can be achieved by, instead of measuring and tracking the injected noise alone, measuring the state difference between the clean and noisy forward passes of the whole network. Note that this is a significant departure from traditional node perturbation in which access to the underlying noise source is required. Concretely, taking the definition of the forward pass given by NP in Eq. 2, we define
@@
where $N=\sum_{l=0}^{L}N_{l}$ is the number of units in the network, $\delta{\mathbf{a}_{l}}=\mathbf{\tilde{a}}_{l}-\mathbf{{a}}_{l}$ is the activity difference between a noisy and a clean pass in layer $l$ and $\delta{\mathbf{a}}=(\delta{\mathbf{a}_{1}},\ldots,\delta{\mathbf{a}_{L}})$ is the concatenation of activity differences. The resulting update rule is referred to as activity-based node perturbation (ANP). Appendix C provides a derivation of this rule.
Here, we have now updated multiple aspects of the NP rule. First, rather than using the measure of noise injected at each layer, we instead measure the total change in activation between the clean and noisy networks. However, relative to BP our derived ANP rule has a biased estimation in gradient measurement. Appendix D describes this bias and shows that, when considering finite noise injection, this bias is measurable in closed form and is interpretable.
Second, this direct use of the activity difference also requires a recomputation of the scale of the perturbation vector. Here we carry the rescaling out by a normalization based upon the activity-difference length and then upscale the signal of this perturbation based upon the network size.
Note that the ANP rule is now a hybridization of the directional derivative and traditional node perturbation. In the limit where the nodes are not interdependent (i.e. when we have a single layer), it is equivalent to INP and upon averaging across many samples can approximate the true gradient with a distinct bias based upon correlations in propagated noise. When used in a layered network structure, it ‘assumes’ that all nodes are mutually independent even though they are not. This is the only remaining approximation for this approach and it is desirable to allow whole network learning without direct access to the noise vectors.
2.2 Increasing NP efficiency through decorrelation
@@
Figure 1: A graphical illustration of the computations required for update measurement for backpropagation (left), node perturbation (middle), and decorrelated activity-based node perturbation (right).
@@
Figure 1 illustrates the difference between different learning rules. In Appendix E, Algorithm 1, we describe decorrelation using the decorrelated activity-based update procedure.
@@ 2.3 Experimental validation
To measure the performance of the algorithms proposed, we ran a set of experiments with the CIFAR-10 (Krizhevsky, 2009) and Tiny ImageNet datasets  (Le & Yang, 2015), using fully-connected and convolutional neural networks, specifically aiming to quantify the performance differences between the traditional (NP), layer-wise iterative (INP) and activity-based (ANP) formulations of NP as well as their decorrelated counterparts. These datasets were chosen as they have previously been shown to yield low performance when training networks using regular node perturbation, demonstrating its limitations (Hiratani et al., 2022). All experiments were repeated using three random seeds, after which performance statistics were aggregated. Further experimental details can be found in Appendix F.
@@ 3.3 Multi-layer networks can be trained effectively with node perturbation
Figure 3 (Bottom panel) shows the performance when training a multi-layer fully connected network. We see all decorrelated algorithms convincingly outperforming their non-decorrelated counterparts. We also see (D)INP outperform (D)ANP and (D)NP, almost keeping up with BP. Performance benefits of (D)ANP over (D)NP are not as clear in these results. Note that DNP performs much better in a three-hidden-layer network than in the single-layer networks, indicating that DNP does facilitate multi-layer credit assignment much more than regular NP, which actually performs worse in the deeper network. As an additional demonstration that our results generalize to other datasets, Appendix J reports results on the SARCOS dataset (Vijayakumar & Schaal, 2000).
Appendix  K explores the scaling of DNP, DANP and DINP to networks with 6 and 9 hidden layers as well as a convolutional neural network trained on CIFAR-10.
@@ 3.5 Resampling allows for scaling to larger systems
To further investigate the capabilities of our approach, we apply BP, DNP, DANP and DINP to the more challenging Tiny ImageNet dataset Le & Yang (2015) using a deeper architecture consisting of three convolutional layers and three fully connected layers. The Tiny ImageNet dataset has 200 classes and thus has a larger output space than CIFAR-10. Larger output spaces are known to be challenging for traditional NP-style algorithms (Hiratani et al., 2022). RGB-values for the images, which are ImageNet images downsized to $64\times 64$ pixels, were standardized as a preprocessing step. For network architecture and learning rates, see Appendix F.
@@ 4 Discussion
In this work, we explored several formulations of NP learning and also introduced a layer-wise decorrelation method which strongly outperforms baseline implementations. In our results we show robust speedups in training compared to regular NP, suggesting that our alternative formulations of NP could prove to be competitive with traditional BP in certain contexts. We attribute this increased efficacy to a more efficient credit assignment by virtue of decorrelated input features at every layer, as well as an attenuation of the bias in NP updates caused by their covariance.
@@
When comparing NP, INP and ANP, a number of observations are in order. First, INP yields the best results stemming from its robust relationship to the directional derivative and underlying gradient. The INP method does, however, require running as many forward passes as there are layers in the network (plus one ‘clean’ pass). This is a drawback for this method and should be considered when analyzing its performance, though it should also be considered that it could remain efficient on hardware in which all forward passes are parallelizable. Second, decorrelated variants of NP and ANP both perform well.
Noise-based learning approaches might also be ideally suited for implementation on neuromorphic hardware (Kaspar et al., 2021). First, as demonstrated in Figure 4, our proposed DANP algorithm scales well even when there is no access to a ‘clean’ model pass. Furthermore, DANP does not require access to the noise signal itself, in contrast to NP. This means that such an approach could be ideally suited for implementation in noisy physical devices (Gokmen, 2021) even in case the noise cannot be measured. Even on traditional hardware architectures, forward passes are often easier to optimize than backward passes and are often significantly faster to compute. This can be especially true for neuromorphic computing approaches, where backward passes require automatic differentiation implementations and a separate computational pipeline for backward passes (Zenke & Neftci, 2021). In these cases, a noise-based approach to learning could prove highly efficient.
@@
Though the presented methods for effective noise-based learning show a great deal of promise, there are a number of additional research steps to be taken. The architectures considered are still relatively shallow, and thus an investigation into how well this approach scales for very deep networks would be beneficial. Testing the scalability of these approaches to tasks of greater complexity is also crucial, as are their application to other network architectures such as residual networks, recurrent networks and attention-based models. Our scaling results do demonstrate that by resampling the noise we achieve better and better alignment with the gradient direction, by virtue of the formal correspondence between directional derivatives and our formulations of node perturbation. In this sense, our approach can be viewed as an alternative to forward accumulation, where we exchange memory complexity (maintaining all partial derivatives) for time complexity (repeated forward passes with resampled noise vectors).
@@
Acknowledgments and Disclosure of Funding
This publication is part of the project Dutch Brain Interface Initiative (DBI2) with project number 024.005.022 of the research programme Gravitation which is (partly) financed by the Dutch Research Council (NWO).

### v6

# Node Perturbation Can Effectively Train Multi-Layer Neural Networks

@@ Abstract
Backpropagation (BP) remains the dominant and most successful method for training parameters of deep neural network models. However, BP relies on two computationally distinct phases, does not provide a satisfactory explanation of biological learning, and can be challenging to apply for training of networks with discontinuities or noisy node dynamics. By comparison, node perturbation (NP), also known as activity-perturbed forward gradients, proposes learning by the injection of noise into network activations, and subsequent measurement of the induced loss change. NP relies on two forward (inference) passes, does not make use of network derivatives, and has been proposed as a model for learning in biological systems. However, standard NP is highly data inefficient and can be unstable due to its unguided noise-based search process. In this work, we develop a number of alternative formulations of NP and relate them to the concept of directional derivatives as well as combining them with a decorrelating mechanism for layer-wise inputs. We find that a closer alignment with directional derivatives together with input decorrelation at every layer theoretically and practically enhances performance of NP learning with large improvements in parameter convergence and much higher performance on the test data, approaching that of BP. Furthermore, our novel formulation allows for application to noisy systems in which the noise process itself is inaccessible.
@@ 1 Introduction
Backpropagation (BP) is the workhorse of modern artificial intelligence. It provides an efficient way of performing multi-layer credit assignment, given a differentiable neural network architecture and loss function (Linnainmaa, 1970). Despite BP’s successes, it requires an auto-differentiation framework for the backward assignment of credit, introducing a distinction between a forward, or inference phase, and a backward, or learning phase, increasing algorithmic complexity and impeding implementation in (neuromorphic) hardware (Kaspar et al., 2021; Zenke and Neftci, 2021). Furthermore, BP has long been criticized for its lack of biological detail and plausibility (Grossberg, 1987; Crick, 1989; Lillicrap et al., 2016), with significant concerns again being the two separate forward and backward phases, but also its reliance on gradient propagation and the non-locality of the information required for credit assignment.
Alternative algorithms have been put forth over the years, though their inability to scale and difficulty in achieving levels of performance comparable to BP have held back their use. Many focus either on greedy local learning rules or on random feedback matrices and projections, which are deemed more plausible than symmetric feedback weights (Dellaferrera and Kreiman, 2022; Frenkel et al., 2021; Kohan et al., 2018; Hinton, 2022)
The category of algorithms we focus on in this work are perturbation algorithms, such as node perturbation (NP) (Dembo and Kailath, 1990; Cauwenberghs, 1992), which require no feedback connectivity at all. In NP, node activations are perturbed by a small amount of random noise. Weights are then updated to produce the perturbed activations in proportion to the degree by which the perturbation improved performance. This method requires a measure of the loss function being optimized on a network both before and after the inclusion of noise. Such an approach is appealing because it leverages the same forward pass twice, rather than relying on two computationally distinct phases. It also does not require non-local information other than the global performance signal.
@@
In this work, we put forward three contributions. First, we reframe the process of node perturbation together with the subsequent measurement of the output loss change in terms of directional derivatives. Directional derivatives have been related to perturbation-based learning before in the context of forward gradient learning (Baydin et al., 2022; Ren et al., 2022) but never used to fully explicate a learning rule in this manner. This provides a more solid theoretical foundation for node perturbation on the whole.
Second, we introduce an approximation, referred to as activity-based node perturbation, which is more efficient and has the additional advantage that it can be implemented in noisy systems such as imprecise hardware implementations (Gokmen, 2021) or biological systems (Faisal et al., 2008), where the noise itself is not directly measurable.
Third, we propose to use a decorrelation method, first described by Ahmad et al. (2023), to de-bias activations. Because NP-style methods directly correlate perturbations of unit activities with changes in a reward signal, decorrelation of these unit activities helps to eliminate confounding effects, making credit assignment more straightforward. In addition, as demonstrated by Hiratani et al. (2022), correlations in the input data lead to more bias in NP’s updates by increasing their covariance. By combining decorrelation with the different NP methods, we find that it is possible to achieve orders of magnitude increase in model convergence speed, with performance levels rivaling networks trained by BP in certain contexts.
@@ 2.1.1 Iterative node perturbation
In the following, we develop a principled approach to node perturbation based learning.
Our goal is to determine the gradient of the loss with respect to the pre-activations in a layer $l$ by use of perturbations. To this end, we consider the partial derivative of the loss with respect to the pre-activation $\mathbf{a}^{i}_{l}$ of unit $i$ in layer $l$ for all $i$. We define a perturbed state as $\mathbf{\tilde{x}}_{k}(h)=f\left(\mathbf{\tilde{a}}_{k}+h\mathbf{m}_{k}\right)$
with $h$ an arbitrary scalar and binary vectors $\mathbf{m}_{k}=\mathbf{e}_{i}$ if $k=l$ with $\mathbf{e}_{i}$ a standard unit vector and $\mathbf{m}_{k}=\mathbf{0}$ otherwise. We may now define the partial derivatives as
@@
An alternative approach is to define perturbations in terms of directional derivatives. Directional derivatives measure the derivative of a function based upon an arbitrary vector direction in its dependent variables.
At this stage a choice can be made: either one may measure this directional derivative for a single layer or for an entire network. Measuring the directional derivative with respect to a specific layer via a perturbation given by
@@
2.1.2 Node perturbation
Further zooming out, one may suppose that they wish to perturb all nodes in an entire network, rather even than just one layer at a time. In regular node perturbation, noise is injected into all layer’s pre-activations and weights are updated in the direction of the noise if the loss improves and in the opposite direction if it worsens. As with iterative node perturbation, two forward passes are required: one clean and one noise-perturbed. During the noisy pass, noise is injected into the pre-activation of each layer to yield a perturbed output
As previously, one can measure a loss differential for an NP-based update. Supposing that the loss $\mathcal{L}$ is measured using the network outputs, the loss difference between the clean and noisy network is given by $\delta{\mathcal{L}}=\mathcal{L}(\mathbf{\tilde{x}}_{L})-\mathcal{L}(\mathbf{x}_{L})$, where $\delta{\mathcal{L}}$ is a scalar measure of the difference in loss induced by the addition of noise to the network. Given this loss difference and the network’s perturbed and unperturbed outputs, we compute a layer-wise learning signal by replacing the gradient $\mathbf{g}_{l}$ in Eq. 1 with a term consisting of the loss difference and a normalized noise vector:
$\Delta\mathbf{W}_{l}^{\textrm{NP}}=\delta{\mathcal{L}}\frac{{\bf\it\epsilon}_{l}}{\sigma^{2}}\mathbf{x}_{l-1}^{\top}\,$ (5)
This update is similar to that of INP, with one key difference, that the normalization term, $\frac{N_{l}}{||v||^{2}_{2}}$, is approximated as the inverse of the node-wise variance, $\frac{1}{\sigma^{2}}$. Thus, regular NP is a network-wide measure of the directional derivative with an approximation of the required normalization.
NP thus explores in a larger parameter space than INP. Importantly, however, it requires knowledge of each layer’s noise vector, $\epsilon_{l}$. This knowledge is hard to attain in real-world cases of noise injection within a neural network as this noise has to be distinguishable from the noise induced by the previous layers.
@@ 2.1.3 Activity-based node perturbation
To improve upon NP and to mitigating the need for exact and separate measurement of the noise at every layer, we propose to approximate the directional derivative across the whole network simultaneously based upon the activity-differences. This can be achieved by, instead of measuring and tracking the injected noise alone, measuring the state difference between the clean and noisy forward passes of the network.
Concretely, taking the definition of the forward pass given by NP in Eq. 4, we define
@@
where $N=\sum_{l=0}^{L}N_{l}$ is the number of units in the network, $\delta{\mathbf{a}_{l}}=\mathbf{\tilde{a}}_{l}-\mathbf{{a}}_{l}$ is the activity difference between a noisy and a clean pass in layer $l$ and $\delta{\mathbf{a}}=(\delta{\mathbf{a}_{1}},\ldots,\delta{\mathbf{a}_{L}})$ is the concatenation of activity differences.
The resulting update rule is referred to as activity-based node perturbation (ANP). Appendix C provides a derivation of this rule.
Here, we have now updated multiple aspects of the NP rule. First, rather than using the measure of isolated noise injected at each layer, we instead measure the change in activation between the clean and noisy networks. This has a theoretical benefit as well as a theoretical downside. As a theoretical benefit, by using the total noise at any given layer, this subsumes noise in all past layers and allows one to update a layer as if there was zero noise added to the preceeding layers. The only correlative noise issues remaining are then with the later network layers. ANP is thus demonstrably better than NP for layers closer to the output of a network.
On the other hand, the noise at each node in a layer is no longer independent and could now have a correlated structure. This is a downside in that it can skew the gradient descent direction, though can be shown to still converge to a minimum in loss. Both of these implications are expanded upon in Appendix C.
Second, this direct use of the activity difference also requires a recomputation of the scale of the perturbation vector and the requisite normalization. Unlike in the case of NP, one cannot assume the size of this normalization.
Note that the ANP rule is now a hybridization of iterative node perturbation and traditional node perturbation. In the limit where the nodes are not interdependent (i.e. when we have a single layer), ANP is equivalent to INP. When used in a layered network structure, ANP allows computation of updates without explicit access to the noise vectors and updates an entire network in the same style as traditional NP.
2.2 Increasing perturbation learning efficiency through decorrelation
@@
In Appendix E, Algorithm 1, we describe decorrelation using the decorrelated activity-based update procedure.
@@ 2.3 Experimental validation
To measure the performance of the algorithms proposed, we ran a set of experiments with the CIFAR-10 (Krizhevsky, 2009) and Tiny ImageNet datasets  (Le and Yang, 2015), using fully-connected and convolutional neural networks, specifically aiming to quantify the performance differences between the traditional (NP), layer-wise iterative (INP) and activity-based (ANP) formulations of NP as well as their decorrelated counterparts. These datasets were chosen as they have previously been shown to yield low performance when training networks using regular node perturbation, demonstrating its limitations (Hiratani et al., 2022). All experiments were repeated using three random seeds, after which performance statistics were aggregated. Further experimental details can be found in Appendix F.
@@ 3.3 Multi-layer networks can be trained effectively with node perturbation
Figure 2 (Bottom panel) shows the performance when training a multi-layer fully connected network. We see all decorrelated algorithms convincingly outperforming their non-decorrelated counterparts. We also see (D)INP outperform (D)ANP and (D)NP, almost keeping up with BP. Performance benefits of (D)ANP over (D)NP are not as clear in these results. Note that DNP performs much better in a three-hidden-layer network than in the single-layer networks, indicating that DNP does facilitate multi-layer credit assignment much more than regular NP, which actually performs worse in the deeper network. As an additional demonstration that our results generalize to other datasets, Appendix J reports results on the SARCOS dataset (Vijayakumar and Schaal, 2000). Appendix  K explores the scaling of DNP, DANP and DINP to networks with 6 and 9 hidden layers as well as a convolutional neural network trained on CIFAR-10.
@@ 3.5 Resampling allows for scaling to larger systems
To further investigate the capabilities of our approach, we apply BP, DNP, DANP and DINP to the more challenging Tiny ImageNet dataset Le and Yang (2015) using a deeper architecture consisting of three convolutional layers and three fully connected layers. The Tiny ImageNet dataset has 200 classes and thus has a larger output space than CIFAR-10. Larger output spaces are known to be challenging for traditional NP-style algorithms (Hiratani et al., 2022). RGB-values for the images, which are ImageNet images downsized to $64\times 64$ pixels, were standardized as a preprocessing step. For network architecture and learning rates, see Appendix F.
@@ 4 Discussion
In this work, we explored several formulations of NP learning and also introduced a layer-wise decorrelation method which strongly outperforms baseline implementations. In our results we show robust speedups in training compared to regular NP, suggesting that our alternative formulations of NP could prove to be competitive with traditional BP in certain contexts. We attribute this increased efficacy to a better alignment with theoretically derived directional derivatives and a more efficient credit assignment by virtue of decorrelated input features at every layer.
@@
When comparing NP, INP and ANP, a number of observations are in order. First, INP yields the best results stemming from its robust relationship to the directional derivative and underlying gradient. The INP method does, however, require running as many forward passes as there are layers in the network (plus one ‘clean’ pass). This is a drawback for this method and should be considered when analyzing its performance, though it should also be considered that it could remain efficient on hardware in which all forward passes are parallelizable. Second, ANP consistently outperforms NP. This can be attributed to its ability to subsume all noise from previous layers into a given layer’s update and to not have any issues of correlation with these past layers. Finally, decorrelation aids all NP variants to achieve higher performance and faster convergence.
Noise-based learning approaches might also be ideally suited for implementation on neuromorphic hardware (Kaspar et al., 2021). First, as demonstrated in Figure 3, our proposed DANP algorithm scales well even when there is no access to a ‘clean’ model pass. Furthermore, DANP does not require access to the noise signal itself, in contrast to NP. This means that such an approach could be ideally suited for implementation in noisy physical devices (Gokmen, 2021) even in case the noise cannot be measured. Even on traditional hardware architectures, forward passes are often easier to optimize than backward passes and are often significantly faster to compute. This can be especially true for neuromorphic computing approaches, where backward passes require automatic differentiation implementations and a separate computational pipeline for backward passes (Zenke and Neftci, 2021). In these cases, a noise-based approach to learning could prove highly efficient.
@@
Though the presented methods for effective noise-based learning show a great deal of promise, there are a number of additional research steps to be taken. Testing the scalability of these approaches to tasks of greater complexity is important, as are their application to other network architectures such as residual networks, recurrent networks and attention-based models. Training accuracy, of even the best NP variants, also show some lag behind backpropagation’s accuracy achievements in deeper networks. Further investigation is needed to determine how to bridge this gap and if it has implications for harder tasks.
Our scaling results do demonstrate that by resampling the noise we achieve better and better alignment with the gradient direction, by virtue of the formal correspondence between directional derivatives and our formulations of node perturbation. In this sense, our approach can be viewed as an alternative to forward accumulation, where we exchange memory complexity (maintaining all partial derivatives) for time complexity (repeated forward passes with resampled noise vectors).